OpenStudy (anonymous):

For the functions  f(x)=5x−8  and  g(x)=3x+5,  find the following. (a)  (f+g)(x) (b)  (f−g)(x)

OpenStudy (anonymous):

@epoweritheta

OpenStudy (anonymous):

a. (f+g)(x) = f(x) +g(x) = 5x-8 +3x+5 8x-3

OpenStudy (anonymous):

lts 3x+5 not -5

OpenStudy (anonymous):

yes. thanks

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

b. (f-g)(x) = f(x)-g(x) 5x-8-3x+5 2x-3 says 2x-3 is wrong. what did I do wrong? @GIL.ojei

OpenStudy (anonymous):

answer : 2x-13 suppose to change the 5 to -5.

OpenStudy (anonymous):

(5x-8)-(3x+5) remember to open up the bracket with minus 5x-8-3x-5= 5x-3x-8-5= 2x-13

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yex. I forgot that. thanks (: